Let x be an element of 0, [0, π]. We consider the inequalities {sinx>=[(2)^1/2]/2, 0<=cosx<(1/2). Which of the following statements is correct?


A. X is an element belonging to [π/4, π/3[;

B. X is an element belonging to ]π/4, 3π/4[

C. X is an element belonging to ]π/3, π]

D. X is an element belonging to ]π/3, π/2]


1
Expert's answer
2021-07-19T14:34:40-0400

Let us consider the system of inequalities "\begin{cases} \sin x\ge\frac{\sqrt{2}}{2},\\ 0\le \cos x<\frac{1}{2}\end{cases}."


The solutions of the inequality "\sin x\ge\frac{\sqrt{2}}{2}" on the interval "[0,\pi]" are all "x\in [\frac{\pi}{4},\frac{3\pi}{4}]." The solutions of the inequality "0\le \cos x<\frac{1}{2}" on the interval "[0,\pi]" are all "x\in ]\frac{\pi}{3},\frac{\pi}{2}]." Taking into account that "[\frac{\pi}{4},\frac{3\pi}{4}]\cap]\frac{\pi}{3},\frac{\pi}{2}]=]\frac{\pi}{3},\frac{\pi}{2}]," we conclude that the solutions of the system are all elements "x" belonging to "]\frac{\pi}{3},\frac{\pi}{2}]."


Answer: D